The Symbol Font wasy

Roland Waldi Institut für Experimentelle Kernphysik Universität Karlsruhe Physikhochhaus, P.O.Box 6980 D-7500 Karlsruhe, Fed. Rep. Germany Version 2.0 – September 1992

The font wasy contains all lasy characters, and a lot more symbols. New characters were modified from the mf files of the standard TEX fonts, and many were designed from scratch. Metafont sources for 5–10pt and a bold 10pt font are available.

An extension to PLAIN-TEX for using the fonts is included in the file wasyfont.tex. This can probably be used in L^{A}TEX documents, but a new L^{A}TEX format with the bindings already included and with wasy replacing the lasy font would be the superiour solution. This version includes all lasy characters at the proper codes (causing some incompatibilities with version 1 of wasy) to make such a procedure easy. The file wasyfont.2 contains substitutes for some macros of wasyfont.tex to be used at installations, that do not support the wasy fonts.

A list of characters with their bindings in wasyfont.tex follows. Some macros are actually compositions of several characters useful in the given context. Macros using symbols which are already available from standard TEX fonts are also included; these are marked with ^{∗}.

general symbols

\male  ♂  \female  ♀
\currency ¤  \cent  ¢
\lozenge ⌑  \kreuz  ✠
\smiley ☺  \blacksmiley ☻
\frownie ☹  \sun  ☼
\checked ✓  \brokenvert ¦
\diameter ⌀  \invdiameter ⍉
\phone  ☎  \recorder  ⌕
\clock  🕒  \permil  ‰
\bell  🔔  \ataribox
\pointer ⇨  \lightning ↯
\agemO  ℧

diagrams

\photon ∼∼∼∼  \gluon

music notes

\eighthnote ♪  \quarternote ♩

\halfnote 𝅗𝅥  \fullnote  𝅝
\twonotes ♫

electrical engineering

\AC ∼  \HF ∼ ∼
\VHF ≋

APL

\APLup  △  \APLdown  ▽
\APLbox  ⎕  \APLinv            ÷ ⎕
\APLleftarrowbox ⍇  \APLrightarrowbox ⍈
\APLuparrowbox  ⍐  \APLdownarrowbox ⍗
\APLinput  ⍞  \APLminus^{∗}  −
\APLlog  ⍟  \APLstar  ★
\APLvert^{∗}  |  \APLvert{\APLdown} |▽
\APLnot^{∗}  ∼  \APLnot{\APLdown}  ∼ ▽
\APLnot{\land}    ∼ ∧  \APLnot{\lor}      ∼ ∨
\APLcirc^{∗}  ∘  \APLcirc{\bot}     ∘ ⊥
\notbackslash^{∗}  \ −  \notslash^{∗}  / −
\APLcomment  ⍝

astronomy

\ascnode ☊  \descnode ☋
\vernal  ♈  \astrosun^{∗} ⊙
\newmoon ●  \fullmoon ◯
\leftmoon ☾  \rightmoon ☽
\mercury ☿  \venus  ♀
\mars  ♂  \jupiter  ♃
\saturn  ♄  \uranus  ⛢
\neptune ♆  \pluto  ♇
\earth  ♁

astrological symbols and zodiacal symbols

\conjunction ☌  \opposition ☍
\aries  ♈  \libra  ♎
\taurus  ♉  \scorpio  ♏
\gemini  ♊  \sagittarius ♐
\cancer  ♋  \capricornus ♑
\leo  ☊  \aquarius  ♒
\virgo  ♍  \pisces  ♓

geometrical shapes

\hexstar  ✶  \varhexstar ✶
\davidsstar ✡  \APLstar  ★
\Circle  ◯  \CIRCLE  ●
\Leftcircle       \LEFTCIRCLE ◖
\Rightcircle      \RIGHTCIRCLE ◗
\LEFTcircle  ◖ ◯  \RIGHTcircle  ◗ ◯
\LEFTarrow ◄  \RIGHTarrow ►
\UParrow  ▲  \DOWNarrow  ▼
\Box  □  \APLbox  ⎕
\XBox  ⊠  \Bowtie  ⨝
\Diamond  ◇  \octagon
\hexagon  ⎔  \varhexagon ⬡
\pentagon  ⬠

general math & physics

\varangle ∢  $\invneg$ ⌐
\leftturn ↺  \rightturn ↻
\diameter ⌀  \therefore ∴

math operators

$a\ocircle b  a ◯ b  $a\logof b  a ⍟ b
$a\oplus^{∗} b  a ⊕ b  $a\otimes^{∗} b  a ⊗ b
$a\le^{∗} b  a ≤ b  $a\ge^{∗} b  a ≥ b
$a\apprle  b  a ≲ b  $a\apprge b  a ≳ b
$a\lhd  b  a ⊲ b  $a\rhd  b  a ⊳ b
$a\unlhd  b  a ⊴ b  $a\unrhd b  a ⊵ b
$a\LHD  b  a ◄ b  $a\RHD  b  a ► b
$a\sqsubset b  a ⊏ b  $a\sqsupset b  a ⊐ b
$a\sqsubseteq^{∗} b a ⊑ b  $a\sqsupseteq^{∗} b a ⊒ b
$a\propto^{∗} b  a ∝ b  $a\varpropto  b a ∝ b
$a\leadsto  b  a ↝ b

integrals (text style)

$\varint a^b f(x)dx ∫_{a}^{b} f(x)dx  $\iint a^b f(x)dx  ∬_{a}^{b} f(x)dx
$\iiint a^b f(x)dx ∭_{a}^{b} f(x)dx  $\varoint a^b f(x)dx ∮_{a}^{b} f(x)dx
$\oiint a^b f(x)dx ∯_{a}^{b} f(x)dx

integrals (display style)

∫ ∬ ∭ ∮ ∯

With the control sequence \newpropto you can change the proportional sign to the thin wasy symbol (∝), which is more distinct from alpha (α) then the default symbol (∝). With the control sequence \newint you can change the TEX integrals from ∫ , ∮ to the vertical ones ∫ , ∮, in display:

∫_{a}^{b} → ∫_{a}^{b}, ∮_{C} → ∮_{C}

There are also a few letters in roman style added (although these and some symbols as ℧, ‰ should be in a separate font, to be created in different styles like italic, sans serif etc. – the wasychr.mf source is prepared for that.

\thorn þ  \Thorn Þ
\dh  ð  \Dh^{∗}  D
\inve ə  \openo ɔ

Examples

“We provide the ♪♫, you provide the ☺” The planets (⊙ → outer space): ☿ ♀ ♁ ♂ asteroids ♃ ♄ ⛢ ♆ ♇. special characters on IBM PC’s: ☺, ☻, ♡, ♢, ♣, ♠, •, ◯, ♂, ♀, ♪, ♫, ☼, ►, ◄, ↕, !!, ¶, §, , \underline{↕}, ↑, ↓, →, ←, ▲, ▼, ¦, ⌐, ¬, … special characters on Atari ST’s: ⇨, , ✓, 🕒, 🔔, ♪, ə, ⌐, ¬, ∮, … tube dimensions: ⌀5 mm, d = 0.5 mm, l = 50 mm display math:

∢(a⃗, b⃗) = 30^{∘}
∏_{x≲5} a_{x} ⍟ b_{x} ≃ ∫_{x≳5} a ◯ b dx  (nonsense.1)
Gauss’ law:  ∭_{V} ∇F(x) d^{3}x = ∯_{S(V)} F(x) da
Stokes’ law:  ∬_{A} [∇ × F(x)] da = ∮_{C(A)} F(x) dl

APL Program: U ← −1 + G ← 2 × ιN ← ⎕ ⍝ generate vectors of odd and even numbers APL keyboard layout:

1 2 3 4 5 6 7 8 9 0 + × ⋄
Q W E R T Y U I O P ← →

A S D F G H J K L [ ]  #
⊢ Z X C V B N M , . /
¨  − < ≤ = ≥ > ≠ ∨ ∧ − ÷ $
? ω ϵ ρ ∼ ↑ ↓ ι ◯ ⋆ { }
α ⌈ ⌊ ▽ △ ∘ ' ⎕ ( ) ∼
⊣ ⊂ ⊃ ∩ ∪ ⊥ ⊤ |  ; : \
I   ∼ ▽ |▽ |△ | ◯ ⍉ ⊖ ⍟ ∼ ∨ ∼ ∧ !  ÷ ⎕ ⍞
\underline{Q} \underline{W} \underline{E} \underline{R} \underline{T} \underline{Y} \underline{U} \underline{I} \underline{O} \underline{P} ⍇ ⍈
\underline{A} \underline{S} \underline{D} \underline{F} \underline{G} \underline{H} \underline{J} \underline{K} \underline{L} ∘ ⊥ ∘ ⊤ ¯,
\underline{△} \underline{Z} \underline{X} \underline{C} \underline{V} \underline{B} \underline{N} \underline{M} ⍝ \ − / −
⊗
⍐ ⍗

simple phonetic notation: corner ['kɔ:nə], this [ðis], thrash [θræ∫] check the appropriate box like this ⊠ or that □ ✓: □ I need the wasy fonts □ I don’t need the wasy fonts

Font Table

wasy:

00 = △ 01 = ⊲ 02 = ⊴ 03 = ⊳ 04 = ⊵ 05 = ∴ 06 = ⌕ 07 = ☎
08 = ✓ 09 = ⇨  0A = 🔔 0B = ♪ 0C = ♩ 0D = 𝅗𝅥 0E = 𝅝 0F = ♫
10 = ◄ 11 = ► 12 = ↯ 13 = ☊ 14 = ☋ 15 = ⍉ 16 = ⍟ 17 = ♈
18 = ⌐ 19 = ♀  1A = ♂ 1B = ¤ 1C = 🕒 1D = ∝ 1E = ∢ 1F = ⌀
20 = ● 21 = ↻ 22 = ↺ 23 = ◯ 24 = ☾ 25 = ☽ 26 = ♁ 27 = ☿
28 =     29 =      2A =     2B =     2C = ☺ 2D = ☻ 2E = ☼ 2F = ☹
30 = ℧ 31 = ⨝ 32 = □ 33 = ◇ 34 = ⊠ 35 = ⌑ 36 = ✠ 37 = ⎔
38 =     39 = ⬡ 3A = ∼ 3B = ↝ 3C = ⊏ 3D = ⊐ 3E = ≲ 3F = ≳
40 = ≋ 41 = ✶ 42 = ✶ 43 = ✡ 44 = ⬠ 45 = ★ 46 = ▽ 47 = ◖
48 = ◗ 49 =      4A =     4B = ▲ 4C = ▼ 4D =     4E =    4F =
50 =     51 =      52 =     53 =     54 =     55 = ə 56 = ☌ 57 = ☍
58 = ♃ 59 = ♄ 5A = ⛢ 5B = ♆ 5C = ♇ 5D = ♉ 5E = ♊ 5F = ♋
60 = ♍ 61 = ♎ 62 = ♏ 63 = ♐ 64 = ♑ 65 = ♒ 66 = ♓ 67 = ¢
68 = ‰ 69 = þ  6A = Þ 6B = ð 6C = ɔ 6D =     6E = ⍐ 6F = ⍗
70 = ⍇ 71 = ⍈ 72 = ∫ 73 = ∬ 74 = ∭ 75 = ∮ 76 = ∯ 77 = ∫
78 = ∬ 79 = ∭ 7A = ∮ 7B = ∯ 7C = ¦ 7D = ⍞ 7E = ⎕ 7F = ⍝

wasyb:

00 = △ 01 = ⊲  02 = ⊴ 03 = ⊳ 04 = ⊵ 05 = ∴ 06 = ⌕ 07 = ☎
08 = ✓ 09 = ⇨  0A = 🔔 0B = ♪ 0C = ♩  0D = 𝅗𝅥 0E = 𝅝 0F = ♫
10 = ◄ 11 = ►  12 = ↯ 13 = ☊ 14 = ☋ 15 = ⍉ 16 = ⍟ 17 = ♈
18 = ⌐ 19 = ♀  1A = ♂ 1B = ¤ 1C = 🕒 1D = ∝ 1E = ∢ 1F = ⌀
20 = ● 21 = ↻ 22 = ↺ 23 = ◯ 24 = ☾ 25 = ☽ 26 = ♁ 27 = ☿
28 =     29 =      2A =     2B =     2C = ☺ 2D = ☻ 2E = ☼ 2F = ☹
30 = ℧ 31 = ⨝  32 = □ 33 = ◇ 34 = ⊠ 35 = ⌑ 36 = ✠ 37 = ⎔
38 =     39 = ⬡  3A = ∼ 3B = ↝ 3C = ⊏ 3D = ⊐ 3E = ≲ 3F = ≳
40 = ≋ 41 = ✶  42 = ✶ 43 = ✡ 44 = ⬠ 45 = ★ 46 = ▽ 47 = ◖
48 = ◗ 49 =      4A =     4B = ▲ 4C = ▼ 4D =     4E =     4F =
50 =     51 =      52 =     53 =     54 =     55 = ə 56 = ☌ 57 = ☍
58 = ♃ 59 = ♄  5A = ⛢ 5B = ♆ 5C = ♇ 5D = ♉ 5E = ♊ 5F = ♋
60 = ♍ 61 = ♎ 62 = ♏ 63 = ♐ 64 = ♑ 65 = ♒ 66 = ♓ 67 = ¢
68 = ‰ 69 = þ  6A = Þ 6B = ð 6C = ɔ 6D =     6E = ⍐ 6F = ⍗
70 = ⍇ 71 = ⍈  72 = ∫ 73 = ∬ 74 = ∭ 75 = ∮ 76 = ∯ 77 = ∫
78 = ∬ 79 = ∭ 7A = ∮ 7B = ∯ 7C = ¦  7D = ⍞ 7E = ⎕ 7F = ⍝

Changes since version 1.0 version 1.1: \varangle has been centered at the math axis version 2.0: new: letters Þ,þ,ð,ə,ɔ,℧ new astrological and zodiacal symbols new symbols permil, cent, ataribox now the full set of lasy is included; for this purpose 9 characters (☺, ☻, ☹, ☼, ♃, ♄, ⛢, ♆, ♇) have changed code! wasyb10 font for bold math added